The original expression is 4x + y = 16
the x = -1
so the first step is to plug in the -1 which would be 4 * -1 + y = 16
Now my question is, how do you solve for y?

Answer by Alan3354(69443)

The original expression is 4x + y = 16
the x = -1
so the first step is to plug in the -1 which would be 4 * -1 + y = 16
Now my question is, how do you solve for y?
==========================
Is x = -1 given? If so:
4x + y = 16
4*(-1) + y = 16
-4 + y = 16
y = 20